ELECTRONIC STRUCTURE CALCULATION OF U(BH_{4})_{4} WITH A CORRELATION CONSISTENT BASIS SET

\underline{ZHIYONG ZHANG} and RUSSELL M. PITZER, Department of Chemical Physics, The Ohio State University, Columbus, OH43210.

A correlation consistent basis set, through g functions, is developed for uranium. Calculations are carried out for U(BH_{4})_{4} using a relativistic effective core potential and the spin-orbit effect included at the configuration interation level. The calculated electronic energy levels are compared with experimental data and previous assignments and calculations. The effect of basis set size on f→f excitation energies is discussed.